Ok I have worked this several times and my answer keeps coming up with six but I don't have that number available on answer sheet what am I doing wrong the problem is x/3 = 2x + 7 what is the number. Equal to the sum
Found 2 solutions by TimothyLamb, rothauserc:
Answer by TimothyLamb(4379)

x/3 = 2x + 7
3(x/3 = 2x + 7)
x = 6x + 21
-5x = 21
x = -21/5
x = -4.2
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by rothauserc(4718)

x/3 = 2x + 7
multiply both sides of = by 3
x = 6x +21
subtract 6x from both sides of =
-5x = 21
x = -21/5